# Learning to Generate Parameters of ConvNets for Unseen Image Data

### v1

@@ V Conclusion and Limitations
In this paper, we found there are correlations among image datasets and the corresponding parameters of a given ConvNet, and explored a new training paradigm for ConvNets. We proposed a new hypernetwork, PudNet, which can directly predict the network parameters for an image unseen dataset with only a single forward propagation. In addition, we attempted to capture the relations among parameters of different layers in a network by a series of adaptive hyper recurrent units. Extensive experimental results demonstrated the effectiveness and efficiency of our method. However, current work primarily focused on exploring the generation of parameters for CNN-based networks. A potential avenue for future research could involve extending parameter generation to other network architectures, e.g., large language models.

### v3

@@ V Conclusion and Future Works
In this paper, we identify correlations among image datasets and the corresponding parameters of a given ConvNet and explore a new training paradigm for ConvNets. We propose a new hypernetwork, called PudNet, which could directly predict the network parameters for an image unseen dataset with only a single forward propagation. In addition, we attempt to capture the relations among parameters across different network layers through a series of adaptive hyper-recurrent units. Extensive experimental results demonstrate the effectiveness and efficiency of our method in both intra-dataset prediction settings and inter-dataset prediction settings.
Nevertheless, the current work primarily delves into parameter generation for target networks based on CNNs. A potential avenue for future research could involve extending parameter generation to large language models (LLMs) [72] or vision language models (VLMs) [73]. For example, LLMs or VLMs commonly employ adapters [74] or LoRA methods [75] to enhance generalization to downstream tasks. This typically requires a certain amount of time for training. In future work, we can explore the potential of our approach in predicting parameters for adapters or LoRA methods, offering a pathway to save training time.